# On the Zariski density of rational curves on IHS manifolds

### v2

@@ Abstract.
In analogy with recent works on $K3$ surfaces, we study the existence of infinitely many ruled divisors on projective irreducible holomorphic symplectic (IHS) manifolds. We prove such an existence result for any projective IHS manifold of $K3^{[n]}$ or generalized Kummer type which is not a variety defined over $\overline{{\mathbb{Q}}}$ with Picard number one. The result is obtained as a combination of the regeneration principle and of a generalization to higher dimension of a controlled degeneration technique, invented by Chen, Gounelas and Liedtke in dimension 2.
@@ 1. Introduction
    All manifolds in $\mathcal{M}$ with Picard number at least two contain infinitely many distinct ample ruled divisors.
@@
    If $[X]\in\mathcal{M}$ corresponds an IHS manifold which is not defined over $\overline{{\mathbb{Q}}}$, then X contains infinitely many distinct ample ruled divisors.
@@
Acknowledgements. We thank F. Charles for bringing [MP12] to our attention, S. Floccari for pointing out Proposition 5.1, S. Tayou and K. O’Grady for valuable comments.
@@ 2. Controlled degenerations
(2.1) ${{\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${E\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{S}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${U}$${\{0\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
    $f:{\mathcal{C}}/V\xrightarrow{\ \ }{\mathcal{S}}$ is a family of stable maps of genus 00 to ${\mathcal{S}}$ over $V$ whose very general fibres ${\mathcal{C}}_{b}$ are smooth for $b\in V$.
@@
and there a a point $e\in E$ and two distinct irreducible components $G_{1}$ and $G_{2}$ of $Y_{e}$ with $f_{*}G_{i}\not=0$ for $i=1,2$. Then there exists an integral curve $F\subset V$ such that $\varphi_{*}F\not=0$, $E\cap F\not=\emptyset$ and ${\mathcal{C}}_{p}$ is singular for all $p\in F$.
Proof of the Controlled degeneration principle 1.7.
@@
be a global section of Hodge type $(2n-1,2n-1)$ of $R^{4n-2}f_{*}{\mathbb{Z}}$ such that $\alpha_{b}$ is the class of primitive curves ruling $D_{b}$. By [CMP24, Proposition 3.1] these curves deform over $U$ to yield, for all $u\in U$ a ruled divisor $D_{u}\in|m{\mathcal{L}}_{u}|$. Up to a base-change we may and will assume that for a very general $u\in U$ and a very general point of $D_{u}$ there exists a unique irreducible rational curve $R_{u}\subset{\mathcal{X}}_{u}$ in the family ruling $D_{u}$. Without loss of generality, we can suppose that the very general set of $U$ where we are taking a very general point of $D_{u}$ contains the central point 00, i.e. we are choosing a very general point of $D_{0}$ and one of the rational curves passing through it. Notice that we cannot suppose that this curve is irreducible, as we obtain it by a flat limit. We then consider the family ${\mathcal{R}}\subset{\mathcal{X}}$ of such rational curves, which by construction contains the central point 00. By eventually applying stable reduction, cf. [CGL22a, Proposition 8.2], we then obtain the following diagram
(2.4) ${{\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${E\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\ \ \ \ \ \ {\mathcal{R}}\subset{\mathcal{X}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\ \ \ \ \ \pi}$${U}$${\{0\}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@
    $f:{\mathcal{C}}/V\xrightarrow{\ \ }{\mathcal{R}}\subset{\mathcal{X}}$ is a flat family of stable maps of genus 00 to ${\mathcal{R}}\subset{\mathcal{X}}$ over $V$;
@@
for all $e\in E$, where ${\mathcal{C}}_{E}:={\mathcal{C}}\times_{V}E$. Indeed if (2.5) did not hold we would have a rationally chain connected surface passing through the very general point of a ruled divisor inside the IHS manifold $\mathcal{X}_{0}$, thus contradicting [Voi16, Theorem 1.3, item (ii)]. Notice that (2.5) is crucial as it corresponds to hypothesis (2.2) in Proposition 2.1. It is automatic in the $2$-dimensional case, as a (complex) $K3$ surface cannot be covered by rational curves, while it requires the finer analysis done in [Voi16] in higher dimensions.
@@
By stability, if this were not the case we could write ${\mathcal{C}}_{e}=G_{1}+G_{2}+G_{more}$, where $G_{1},G_{2}$ are two irreducible components of ${\mathcal{C}}_{e}$ such that $f_{*}(G_{i})\not=0,\ i=1,2$. We could then apply Proposition 2.1 with $Y={\mathcal{C}}_{E}$ and get an integral curve $F\subset V$ such that $\varphi_{*}F\not=0$, $E\cap F\not=\emptyset$ and ${\mathcal{C}}_{p}$ is singular for all $p\in F$. Therefore let us write ${\mathcal{C}}_{p}=M_{p}+N_{p}$ for all $p\in F$ where $M_{p}$ and $N_{p}$ are union of irreducible components of ${\mathcal{C}}_{p}$ with $f_{*}M_{p}\not=0\not=f_{*}N_{p}$. In particular, for $p\in E\cap F$ we would have $[f_{*}M_{p}]+[f_{*}N_{p}]=\alpha_{0}$. Here we denoted by $[f_{*}M_{p}]$ and $[f_{*}N_{p}]$ the cohomology classes of $1$-cycles in $\mathcal{X}_{p}$ and $\alpha_{0}$ is as in (2.3). As $\alpha_{0}$ is primitive, the two components $[f_{*}M_{p}],\,[f_{*}N_{p}]$ generate a rank two ${\mathbb{Z}}$-module. By hypothesis, the dual of $\Lambda$ is given by algebraic classes of type $(2n-1,2n-1)$ on $X$ which therefore form a ${\mathbb{Z}}$-module of rank two.
@@
Hence, by item (jj) above $[f_{*}M_{p}]$ and $[f_{*}N_{p}]$ do not belong to $\Sigma^{\vee}$.
@@ 3. Applications
Let $\mathcal{M}$ be a component of the moduli space of $M$ lattice-polarized IHS manifold of $K3^{[n]}$ or generalized Kummer type, where $M$ is an even indefinite lattice of rank 2. Then the locus of points of $\mathcal{M}$ which are birational to a Hilbert scheme of points on a K3 surface or to a generalized Kummer manifold is dense.
@@
We prove non-emptyness for the $K3^{[n]}$ type, the proof for the generalized Kummer type is identical with the obvious modifications. We denote $\Lambda=U^{\oplus 3}\oplus E_{8}(-1)^{\oplus 2}\oplus\mathbb{Z}\ell$ with $\ell^{2}=-2(n-1)$. Observe that any sublattice $R\subset\Lambda$ containing $\ell$ splits as $\mathbb{Z}\ell\oplus\ell^{\perp_{R}}$, by [GHS13, Lemma 7.2] and [Huy16, Chapter 14, (0.2)]. A component $\mathcal{M}$ corresponds to an orbit of isometric embeddings $M\hookrightarrow\Lambda$; we fix one such embedding. If $M\subset\Lambda$ contains $\sigma(\ell)$ for some $\sigma$ in the $M$-polarized monodromy, by the observation above a very general point $p$ in the period domain of $\mathcal{M}$ corresponds to the same Hodge structure of the Hilbert scheme of points on some K3 surface $S$, whose Néron-Severi is isomorphic to $\ell^{\perp_{M}}$. By the surjectivity of the period map for marked IHS manifolds and the Global Torelli Theorem, there is a $M$-polarized IHS, birational to $S^{[n]}$, whose period point is $p$. By [Huy99, Theorem 4.6], we know that birational $M$-polarized IHS correspond to inseparable points in the moduli space. Hence the non-emptyness of IHS birational to $S^{[n]}$ yields the non-emptyness of IHS isomorphic to an Hilbert scheme once we prove that the nef cone of $S^{[n]}$ has non-empty intersection with the copy of $M$ in $\operatorname{Pic}(S^{[n]})$: this is clear, since the two lattices are the same for the very general case.
If instead $\sigma(\ell)\notin M$ for any $\sigma$, we consider the saturated $R$ of $\langle M,\ell\rangle$ and we repeat the argument. Note that $R$ is in the form $N\oplus\mathbb{Z}\ell$, by the same lattice-theoretic argument as above. In this case the Néron-Severi group of $S$ is isometric to $N$; its intersection with the nef cone of $S^{[n]}$ lies then in the wall associated to the Hilbert-Chow contraction, hence any non-zero element in the intersection $M\cap N$, whose rank is at least one, yields a pseudo-polarization of the Hilbert scheme and we are done. ∎
@@
Let $\Lambda$ be a rank two lattice such that $\Lambda\subset\operatorname{Pic}(X)$ and there is an ample class contained in $\Lambda$. Let us consider the moduli space of $\Lambda$-polarized IHS manifold, and let $\mathcal{M}$ be a component that contains $X$ with the above given inclusion $\Lambda\subset\operatorname{Pic}(X)$. Thanks to Lemma 3.1, we can pick a point in the component of the moduli space in which $X$ lies which parametrizes the punctual Hilbert scheme $S^{[n]}$ of a projective $K3$ surface $S$ (resp. the generalized Kummer manifold $K_{n}(A)$ of an abelian variety). We therefore have a family $f:\mathcal{X}\xrightarrow{\ \ }B$ of projective IHS manifolds such that for a point $b_{0}\in B$ we have $X_{0}:=f^{-1}(b_{0})\cong X$, for $b_{1}\in B$ we have $X_{1}:=f^{-1}(b_{1})\cong S^{[n]}$ (resp. $\cong K_{n}(A)$) and $\forall b\in B$ the Picard number of $X_{b}:=f^{-1}(b)$ is at least 2. We know that $S^{[n]}$ contains infinitely many divisors ruled by integral rational curves (by taking infinitely many rational curves in $S$) and ditto for $K_{n}(A)$ (this time by the proof of [MP25, Theorem 1.6, item (2)]). By [MP25, Theorem 1.4] we can regenerate these divisors and get infinitely many ample ruled divisors for very general $b\in B$. Then it suffices to apply the controlled degeneration principle Theorem 1.7 to conclude. ∎
@@
Indeed, as in [Huy16, Chapter 17, Lemma 2.2], a line bundle $M$ on $X_{\eta}$ can be seen as a family of line bundles over $X$, parametrized by a base $\textrm{Spec}(A)$ (where $A$ is an algebra whose quotient field is $k(\eta))$. A section of $M$ corresponding to a ruled divisor on $X_{\eta}$ gives therefore a family of sections over Spec(A) giving a ruled divisor on each fibre $X_{a}\cong X$. However, as ruled divisors on an IHS manifold are rigid in their linear systems, these sections need to be locally constant, hence they correspond to a single ruled divisor on $X$.
@@
Therefore, by Theorem 1.4, item (1), we know that for all points $b\in B_{0}$ there are countably many distinct ruled divisors, and we take a regeneration of them using [MP25, Theorem 1.4]. As we are regenerating divisors which are distinct, their regenerations are also distinct over the generic point $X_{\eta}$. By the above Claim we obtain countably many distinct ruled divisors on $X$. ∎

### v3

@@ Abstract.
In analogy with recent works on $K3$ surfaces, we study the existence of infinitely many ruled divisors on projective irreducible holomorphic symplectic (IHS) manifolds. We prove such an existence result for any projective IHS manifold of $K3^{[n]}$ or generalized Kummer type which is not a variety defined over $\overline{{\mathbb{Q}}}$ with Picard number one or maximal. The result is obtained as a combination of the regeneration principle and of a generalization to higher dimension of a controlled degeneration technique, invented by Chen, Gounelas and Liedtke in dimension 2.
@@ 1. Introduction
    All manifolds in $\mathcal{M}$ with Picard number at least two and strictly smaller than the maximal rank contain infinitely many distinct ample ruled divisors.
@@
    If $X]\in\mathcal{M}$ corresponds an IHS manifold which is not defined over $\overline{{\mathbb{Q}}}$, then X contains infinitely many distinct ample ruled divisors.
Remark 1.5.
In a forthcoming paper [Kir25], Kirschmer proves that there are no reflective lorentzian lattices of rank 21. These are lattices of signature $(1,20)$ whose isometry group is a finite extension of the group generated by reflections. By the proof of the birational Morrison–Kawamata cone conjecture for IHS manifolds [Mar11, Theorem 6.25], it follows that a IHS manifold with finite birational automorphism group has Picard lattice which is reflective lorentzian. Therefore, manifolds of $K3^{[n]}$ type with maximal Picard rank have always infinite birational automorphism groups and they have infinitely many distinct big and nef ruled divisors by [MP25, Theorem 1.7].
@@
Acknowledgements. We thank F. Charles for bringing [MP12] to our attention, S. Floccari for pointing out Proposition 5.1, S. Brandhorst, M. Kirschmer, S. Tayou and K. O’Grady for valuable comments.
@@ 2. Controlled degenerations
[…]
@@
    $f:{\mathcal{C}}/V\xrightarrow{\ \ }{\mathcal{S}}$ is a family of stable maps of genus 0 to $\mathcal{S}$} over $V$ whose very general fibres ${\mathcal{C}}_{b}$ are smooth for $b\in V$.
@@
and there are a point $e\in E$ and two distinct irreducible components $G_{1}$ and $G_{2}$ of $Y_{e}$ with $f_{*}G_{i}\not=0$ for $i=1,2$. Then there exists an integral curve $F\subset V$ such that $\varphi_{*}F\not=0$, $E\cap F\not$$\emptyset$ and ${\mathcal{C}}_{p}$ is singular for all $p\in F$.
Proof of the Controlled degeneration principle 1.8.
@@
be a global section of Hodge type (2n1,2n1)(2n-1,2n-1) of $R^{4n-2}f_{*}{\mathbb{Z}}$ such that $\alpha_{b}$ is the class of primitive curves ruling $D_{b}$. By [CMP24, Proposition 3.1] these curves deform over $U$ to yield, for all $u\in U$ a ruled divisor $D_{u}\in|m{\mathcal{L}}_{u}|$. The locus in $D_{u}$ given by points where a positive dimensional family of rational curves passes is a closed subset of smaller dimension for all $u\in U$ by [Voi16, Theorem 1.3 item (ii)]. Therefore, up to a base change we may and will assume that for any point of $D_{u}$ outside of the closed locus above, there exists a unique rational curve $R_{u}\subset{\mathcal{X}}_{u}$ in the family ruling $D_{u}$. Furthermore, for $u\in U$ very general, this curve is irreducible. Notice that by hypothesis this holds also on very general points of $B$.
We then consider the family ${\mathcal{R}}\subset{\mathcal{X}}$ of such rational curves, which by construction contains the central point 0. By eventually applying stable reduction, cf. [CGL22a, Proposition 8.2], we then obtain the following diagram
[…]
@@
    $f:{\mathcal{C}}/V\xrightarrow{\ \ }{\mathcal{R}}\subset{\mathcal{X}}$ is a flat family of stable maps of genus 0 to ${\mathcal{R}}\subset{\mathcal{X}}$ over $V$;
@@
for all $e\in E$, where ${\mathcal{C}}_{E}:={\mathcal{C}}\times_{V}E$. Indeed if (2.5) did not hold, fix an irreducible component of $f({\mathcal{C}}_{E})$: we would have a rationally chain connected surface passing through the very general point of this irreducible ruled divisor inside the IHS manifold $\mathcal{X}_{0}$, thus contradicting [Voi16, Theorem 1.3, item (ii)]. Notice that (2.5) is crucial as it corresponds to hypothesis (2.2) in Proposition 2.1. It is automatic in the $2$-dimensional case, as a (complex) $K3$ surface cannot be covered by rational curves, while it requires the finer analysis done in [Voi16] in higher dimensions.
@@
By stability, if this were not the case we could write ${\mathcal{C}}_{e}=G_{1}+G_{2}+G_{more}$, where $G_{1},G_{2}$ are two irreducible components of ${\mathcal{C}}_{e}$ such that $f_{*}(G_{i})\not=0,\ i=1,2$. We could then apply Proposition 2.1 with $Y={\mathcal{C}}_{E}$ and get an integral curve $F\subset V$ such that $\varphi_{*}F\not=0$, $E\cap F\not$$\emptyset$ and ${\mathcal{C}}_{p}$ is singular for all $p\in F$. Therefore let us write ${\mathcal{C}}_{p}=M_{p}+N_{p}$ for all $p\in F$ where $M_{p}$ and $N_{p}$ are union of irreducible components of ${\mathcal{C}}_{p}$ with $f_{*}M_{p}\not=0\not=f_{*}N_{p}$. In particular, for $p\in E\cap F$ we would have $f_{*}M_{p}]+[f_{*}N_{p}]=\alpha_{0}$. Here we denoted by $f_{*}M_{p}]$ and $f_{*}N_{p}]$ the cohomology classes of $1$-cycles in $\mathcal{X}_{p}$ and $\alpha_{0}$ is as in (2.3). As $\alpha_{0}$ is primitive, the two components $f_{*}M_{p}],\,[f_{*}N_{p}]$ generate a rank two $\mathbb{Z}$}-module. By hypothesis, the dual of $\Lambda$ is given by algebraic classes of type (2n1,2n1)(2n-1,2n-1) on $X$ which therefore form a $\mathbb{Z}$}-module of rank two.
@@
Hence, by item (jj) above $f_{*}M_{p}]$ and $f_{*}N_{p}]$ do not belong to $\Sigma^{\vee}$.
@@ 3. Applications
Let $\mathcal{M}$ be a component of the moduli space of $\Lambda$ lattice-polarized IHS manifold of $K3^{[n]}$ or generalized Kummer type, where $\Lambda$ is an even indefinite lattice of rank between 2 and $b_{2}-3$. Then the locus of points of $\mathcal{M}$ which are birational to a Hilbert scheme of points on a K3 surface or to a generalized Kummer manifold is dense.
@@
We prove non-emptyness for the $K3^{[n]}$ type, the proof for the generalized Kummer type is identical with the obvious modifications. We denote $L=U^{\oplus 3}\oplus E_{8}(-1)^{\oplus 2}\oplus\mathbb{Z}\ell$ with $\ell^{2}=-2(n-1)$. Observe that any sublattice $R\subset L$ containing $\ell$ splits as $\mathbb{Z}\ell\oplus\ell^{\perp_{R}}$, by [GHS13, Lemma 7.2] and [Huy16, Chapter 14, (0.2)]. A component $\mathcal{M}$ corresponds to an orbit of isometric embeddings $\Lambda\hookrightarrow L$; we fix one such embedding. If $\Lambda\subset L$ contains $\sigma(\ell)$ for some $\sigma$ in the $M$-polarized monodromy, by the observation above a very general point $p$ in the period domain of $\mathcal{M}$ corresponds to the same Hodge structure of the Hilbert scheme of points on some K3 surface $S$, whose Néron-Severi is isomorphic to $\ell^{\perp_{\Lambda}}$. By the surjectivity of the period map for marked IHS manifolds and the Global Torelli Theorem, there is a $\Lambda$-polarized IHS, birational to $S^{[n]}$, whose period point is $p$. By [Huy99, Theorem 4.6], we know that birational $\Lambda$-polarized IHS correspond to inseparable points in the moduli space. Hence the non-emptyness of IHS birational to $S^{[n]}$ yields the non-emptyness of IHS isomorphic to an Hilbert scheme once we prove that the nef cone of $S^{[n]}$ has non-empty intersection with the copy of $\Lambda$ in $\operatorname{Pic}(S^{[n]})$: this is clear, since the two lattices are the same for the very general case.
If instead $\sigma(\ell)\notin\Lambda$ for any $\sigma$, we consider the saturated $R$ of $\langle\Lambda,\ell\rangle$ and we repeat the argument. Note that $R$ is in the form $N\oplus\mathbb{Z}\ell$, by the same lattice-theoretic argument as above. In this case the Néron-Severi group of $S$ is isometric to $N$; its intersection with the nef cone of $S^{[n]}$ lies then in the wall associated to the Hilbert-Chow contraction, hence any non-zero element in the intersection $\Lambda\cap N$, whose rank is at least one, yields a pseudo-polarization of the Hilbert scheme and we are done. ∎
@@
Let $\Lambda$ be a lattice of rank between two and $b_{2}(X)-3$ such that $\Lambda=\operatorname{Pic}(X)$ and there is an ample class contained in $\Lambda$. Let us consider the moduli space of $\Lambda$-polarized IHS manifold, and let $\mathcal{M}$ be a component that contains $X$ with the above given inclusion $\Lambda=\operatorname{Pic}(X)$. Thanks to Lemma 3.1, we can pick a point in the component of the moduli space in which $X$ lies and which parametrizes the punctual Hilbert scheme $S^{[n]}$ of a projective $K3$ surface $S$ (resp. the generalized Kummer manifold $K_{n}(A)$ of an abelian variety). We therefore have a family $f:\mathcal{X}\xrightarrow{\ \ }B$ of projective IHS manifolds such that for a point $b_{0}\in B$ we have $X_{0}:=f^{-1}(b_{0})\cong X$, for $b_{1}\in B$ we have $X_{1}:=f^{-1}(b_{1})\cong S^{[n]}$ (resp. $\cong K_{n}(A)$) and $\forall b\in B$ the Picard number of $X_{b}:=f^{-1}(b)$ is at least 2. We know that $S^{[n]}$ contains infinitely many divisors ruled by integral rational curves (by taking infinitely many rational curves in $S$) and ditto for $K_{n}(A)$ (this time by the proof of [MP25, Theorem 1.6, item (2)]). By [MP25, Theorem 1.4] we can regenerate these divisors and get infinitely many ample ruled divisors for very general $b\in B$. Then it suffices to apply the controlled degeneration principle Theorem 1.8 to conclude. ∎
@@
We have moreover the following:
Lemma 3.3.
Let $B$ a positive dimensional subscheme of a component $\mathcal{M}$ of the moduli space of polarized IHS manifolds with $b_{2}\geq 5$. Let $B_{fin}$ be the subset of $B$ corresponding to points $(X,H)$ such that $X$ has maximal Picard rank and $Bir(X)$ is finite. Then the set $B_{fin}$ is a finite set.
Proof.
By the birational Morrison–Kawamata cone conjecture [Mar11, Theorem 6.25], a IHS manifold with finite birational automorphism group has a reflective lorentzian Picard lattice. By [Nik82, Theorem 5.2.1], in any given rank (at least three), these lattices fall into finitely many isometry classes, up to scaling. As we are restricting to IHS manifolds with $b_{2}\geq 5$, this applies to all lattices $\operatorname{Pic}(X)$ such that $(X,H)\in B_{fin}$ for some $H$. Furthermore, we have that $Bir(X)$ is commensurable with $O(\operatorname{Pic}(X))/W_{Pex}(\operatorname{Pic}(X))$, where $W_{Pex}(\operatorname{Pic}(X))$ is the group generated by reflections on stably prime exceptional divisors by [Mar11, Theorem 6.18, item (5)]. As stably prime exceptional divisors have bounded square in any deformation class as a consequence of [Mar11, Proposition 6.2], for any reflective lorentzian lattice there is a finite number (possibly zero) of rescalings such that it corresponds to the Picard lattice of some $(X,H)\in B_{fin}$. This implies that the natural forgetful projection $B_{fin}\xrightarrow{\ \ }\{\text{IHS manifolds}\}/\sim$ mapping $(X,H)$ to the isomorphism class of $X$ has finite image. However, by [Huy18, Theorem 0.1], the fibers of this forgetful projection are finite, hence the lemma follows. ∎
Notice that, by 1.5, in the case of manifolds of $K3^{[n]}$ type the set $B_{fin}$ is actually empty.
@@
Indeed, as in [Huy16, Chapter 17, Lemma 2.2], a line bundle $M$ on $X_{\eta}$ can be seen as a family of line bundles over $X$, parametrized by a base $\textrm{Spec}(A)$ (where $A$ is an algebra whose quotient field is $k(\eta))$. A section of $M$ giving a ruled divisor on $X_{\eta}$ gives therefore a family of sections over Spec(A) giving a ruled divisor on each fibre $X_{a}\cong X$. However, as ruled divisors on an IHS manifold are rigid in their linear systems, these sections need to be locally constant, hence they correspond to a single ruled divisor on $X$.
@@
Therefore, by Theorem 1.4, item (1), we know that for all points $b\in B_{0}$ that do not have maximal Picard rank there are countably many distinct ruled divisors, and we take a regeneration of them using [MP25, Theorem 1.4]. As we are regenerating divisors which are distinct, their regenerations are also distinct over the generic point $X_{\eta}$. By the above Claim we obtain countably many distinct ruled divisors on $X$.
It remains to prove the theorem in the case when all points of $B_{0}$ have maximal Picard rank. If such a $(X_{b},H_{b})$, $b\in B_{0}$, has infinite birational automorphism group, we can still use the same argument as before, as by [MP25] $X_{b}$ will have countably many distinct uniruled divisors. However, by Lemma 3.3, the subset $B_{fin}$ of $B_{0}$ consisting of pairs $(X_{b},H_{b})$ with maximal Picard rank and finite birational automorphism group is finite, hence there are at least countably many points in $B_{0}\setminus B_{fin}$ and the Theorem follows. ∎